OpenStudy (anonymous):

PLEASE HELP ME!!! Solve the equation: -3 2/5 + b = 8 1/5 A. b=11 3/5 B. b= 4 4/5 C. b= 5 3/5 D. b= 11 1/5

OpenStudy (anonymous):

A. b=11 3/5 because you need to add the left side to the right to isolate the b on the left side.

OpenStudy (anonymous):

8+3 =11 and 1/5+2/5 = 3/5

OpenStudy (anonymous):

thank you so much!!! :D